# arXiv:2602.23780v1[eess.SP] 27 Feb 2026

### 27 Feb 2026 v1

@@ 2.3 Discussion about polynomials in the Euclidean space
Note that a similar result to Definition 1 holds in $\mathbb{R}^{d}$. Indeed, let $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ be even. Then
@@
Fix then any even $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ and $\varepsilon>0$. In either case, observe it is in fact immediate to generalize certain previous results. Indeed, the proof of Theorem 1 is extrapolated as is to $\mathbb{R}^{d}$, from which we get that
@@
$(p*\varphi_{\varepsilon})(x)$ $=\sum_{\alpha\leq n}a_{\alpha}x^{\alpha}+\sum_{\alpha\leq n}\sum_{\begin{subarray}{c}\beta<\alpha\\ |\alpha-\beta|\in 2\mathbb{N}\end{subarray}}\binom{\alpha}{\beta}a_{\alpha}x^{\beta}\varepsilon^{d|\alpha-\beta|}c_{\alpha-\beta}$
$=:p(x)+p_{1}(x),\quad(x\in\mathbb{R}^{d}).$
which is similar to the results of equation (4) in Corollary 1. That is, the convolution of an $n$ degree polynomial in $\mathbb{R}^{d}$ with even Schwartz convolution kernel, results in the sum of the polynomial itself with a polynomial of degree n2n-2. This therefore implies that the results of Corollary 2 can be extrapolated, by noting that $|\alpha|=1\iff\alpha\in\{(1,\dots,0),(0,1\dots,0),\dots,(0,\dots,1)\}$, and therefore $x\mapsto x^{\alpha}=x_{i}$ for some $i\in\{1,\dots,d\}$. Thus, $\int_{\mathbb{R}^{d}}(x-y)^{\alpha}\varphi_{\varepsilon}(y)dy=x_{i}-\int_{\mathbb{R}^{d}}x_{i}\varphi_{\varepsilon}(y)dy=s_{i}$ for some $i\in\{1,\dots,d\}$ and all $x\in\mathbb{R}^{d}$. Thus if $p\in\mathscr{P}_{1}(\mathbb{R}^{d})$ we get
@@
that is, polynomials in $\mathscr{P}_{1}(\mathbb{R}^{d})$ and $\mathscr{P}_{0}(\mathbb{R}^{d})$ are invariant to the convolution with even Schwartz convolution kernel. Thus, we can reach the same conclusions as in Theorem 2 with an identical proof. That is, convolution with an even convolution kernel is an isomorphism of $\mathscr{P}_{n}(\mathbb{R}^{d})$ onto itself for all $n\in\mathbb{N}_{0}$, and we can compute the inverse of the convolution similar to equations (7) and (8). Indeed, if $n$ is even the inverse of $p\in\mathscr{P}_{n}(\mathbb{R}^{d})\mapsto T_{\varepsilon}(p)=p*\varphi_{\varepsilon}\in C^{\infty}(\mathbb{R}^{d})$ with $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ even is
@@ 3 Applications
In this section we present two particular applications of the results in Section 2. First we validate numerically Theorem 2, and then we apply the results of Theorem 3 to an application in the reconstruction of a sinusoidal function.
Figure 1: Representation of a numerical validation of Theorem 2. The left picture represents as a solid black line the original polynomial $p$ which is the $n=50$ order Taylor polynomial of the function $x\in\mathbb{R}\mapsto\sin(5x)+\sin(3x)$. It also represents as a blue solid line its convolution $p*\varphi_{\varepsilon}$ with convolution kernel $s\mapsto\varphi(x)=\operatorname{ind}_{(-1,1)}\exp(-1/(1-x^{2}))$ and parameter $\varepsilon=0.9$. The dashed red line represents $T_{\varepsilon}^{-1}(p*\varphi_{\varepsilon})$ computed using equation (7). The right picture represents the error between the original polynomial and the inverse of the convolution.
3.1 Validation of Theorem 2
In order to numerically validate Theorem 2, we present Figure 1. The polynomial is an $n=50$ Taylor polynomial of the function $x\in\mathbb{R}\mapsto\sin(3x)+\sin(5x)$. Due to the fact that, numerically speaking, the domain of the original polynomial is always bounded, it can be seen that at the ends of the domain the inverse of the convolution and the original polynomial differs the most. This is a fact on how convolution is computed numerically. Due to being the functions actually finite sequences, a zero padding is needed in order to be able to compute the convolution numerically. Then, once computed, we take back the same number of samples as the original sequence in order to compare it with the original function in its domain of definition. Nevertheless, it can be seen that, once the effect of the edges is negligible, that Figure 1 validates numerically Theorem 2, recovering $p$ from $p*\varphi_{\varepsilon}$ by applying Equation (7). Note that the used convolution kernel has compact support and zeros in its Fourier transform (cf. Theorem 3).
@@ 3.2 Applications to Signal Processing
We are going to consider the problem of deconvolution of an arbitrary function. Suppose that the convolution kernel is the Gaussian Kernel, $\varphi\in\mathscr{S}(\mathbb{R})$. Suppose then, that some function $f:\mathbb{R}\to\mathbb{R}$ has been convolved with $\varphi_{\varepsilon}$ for some $\varepsilon>0$. Suppose further that $f$ satisfies any condition of Theorem 3, Then, by letting $n\to\infty$ in Equation (7) we could recover $f$ (in the appropriate convergence).
The traditional approach to deconvolution consists on working on the Fourier transformed space. For example, under the appropriate conditions, for $\varphi$ the Gaussian Kernel and $f\in\mathscr{S}(\mathbb{R})$ it holds that $\widehat{f*\varphi}(\xi)=\widehat{f}\widehat{\varphi}(\xi)\Longrightarrow f=\mathcal{F}^{-1}\left(\widehat{f*\varphi}e^{\operatorname{id}^{2}}\right)$. Note this is ill-posed numerically speaking, since $x\mapsto e^{x^{2}}$ growths exponentially fast, and in particular, this function does not posses inverse Fourier transform. However, by using the results of Theorem 3, we can reconstruct $f$ directly from the samples of $f*\varphi$, without the need of passing to the frequency domain.
Consider then, the function $t\in\mathbb{R}\mapsto f(t)=\sin(5t)+\sin(3t)$ (we denote $t$ for time) as in Figure 1 and let $\varphi$ be the Gaussian convolution kernel. Since $f$ defines a regular tempered distribution, the assumptions of Theorem 3 hold.
Figure 2: Representation of the numerical validation of Theorem 3 under ideal conditions. Left picture represents in black the original function $t\mapsto f(t)=\sin(5t)+\sin(3t)$, in blue its convolution $T_{\varepsilon}(f)$ with the Gaussian Kernel $\varphi$ with $\varepsilon=0.55$, and as dashed red line its reconstruction $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$ using Equation (9) for $n=90$. Right picture represents in black the Fast Fourier Transform (FFT) of $f$, in green the FFT of $\varphi$, in blue the FFT of $T_{\varepsilon}(f)$ and in red the FFT of the reconstructed function $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$.
Note $T_{\varepsilon}(f)=f*\varphi_{\varepsilon}$ for $\varepsilon>0$ is the available data. In Figure 2 we can see the results of applying Theorem 3 to $T_{\varepsilon}(f)$. As it can be seen, the signal reconstruction is almost perfect. Theoretically speaking, we would need to let $n\to\infty$ for (distributional) convergence, but at $n\geq 100$ the numerical errors start to affect the method, and the recovery of the signal for $n=90$ is sufficiently good, numerically speaking, as can be seen from the time and frequency domains. Indeed, by looking at the spectra of $f$, $T_{\varepsilon}(f)$ and $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$, it is well validated that the method works, since we were capable of recovering (numerically speaking) the original function.
@@
Figure 3: Representation of the numerical validation of Theorem 3 under non ideal conditions. Top and bottom left pictures represent in black the original function $t\mapsto f(t)=\sin(5t)+\sin(3t)$, with a blue line $T_{\varepsilon}(f)$ with $\varepsilon=0.55$ and $\varphi$ the Gaussian Kernel, as a yellow line $T_{\varepsilon}(f)+\eta$, where $\eta$ is noise generated using a normal distribution with zero mean and variance 0, and with a red line $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)$ where $n=90$. Top right picture depicts as a black line the FFT of $f$, as a yellow line the FFT of $T_{\varepsilon}(f)+\eta$ and as a red line $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)$, i.e., the application of Equation (9) to $T_{\varepsilon}(f)+\eta$. The bottom right picture shows with a black line the FFT of $f$, with a green line the FFT of the used filter $t\mapsto h(t)=2\operatorname{sinc}(2t)$, with a yellow line the FFT of $T_{\varepsilon}(f)+\eta$ and with a red line the FFT of $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)*h$, that is, $g:=T_{\varepsilon}(f)+\eta$.
@@ 4 Conclusions
In this work we have shown that even Schwartz convolution (i.e., convolution with even Schwartz convolution kernels) is an automorphism between polynomial spaces of finite order. In particular, we obtained an explicit formula for the inversion of the convolution, and numerically validated it with the Taylor polynomial of a sinusoidal function. Moreover, we have shown how these results can be extended to a bigger class of functions and tempered distributions, having a new iterative formula for the inverse of the convolution of these functions and distributions with Schwartz kernels. We have numerically validated this extension with an application to signal processing in which a sinusoidal function has been recovered after convolution, either in a noiseless scenario and in a less than unity signal to noise ratio situation.

### 17 Apr 2026 v2

@@ 2.3 Discussion about polynomials in the Euclidean space
Note that a similar result to Definition 2.1 holds in $\mathbb{R}^{d}$. Indeed, let $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ be even. Then
@@
Fix then any even $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ and $\varepsilon>0$. In either case, observe it is in fact immediate to generalize certain previous results. Indeed, the proof of Theorem 2.3 is extrapolated as is to $\mathbb{R}^{d}$, from which we get that
@@
$(p*\varphi_{\varepsilon})(x)$ $=\sum_{\alpha\leq n}a_{\alpha}x^{\alpha}+\sum_{\alpha\leq n}\sum_{\begin{subarray}{c}\beta<\alpha\\ |\alpha-\beta|\in 2\mathbb{N}\end{subarray}}\binom{\alpha}{\beta}a_{\alpha}x^{\beta}\varepsilon^{d|\alpha-\beta|}c_{\alpha-\beta}=:p(x)+p_{1}(x),\quad(x\in\mathbb{R}^{d}).$
which is similar to the results of equation (4) in Corollary 2.5. That is, the convolution of an $n$ degree polynomial in $\mathbb{R}^{d}$ with even Schwartz convolution kernel, results in the sum of the polynomial itself with a polynomial of degree n2n-2. This therefore implies that the results of Corollary 2.8 can be extrapolated, by noting that $|\alpha|=1\iff\alpha\in\{(1,\dots,0),(0,1\dots,0),\dots,(0,\dots,1)\}$, and therefore $x\mapsto x^{\alpha}=x_{i}$ for some $i\in\{1,\dots,d\}$. Thus, $\int_{\mathbb{R}^{d}}(x-y)^{\alpha}\varphi_{\varepsilon}(y)dy=x_{i}-\int_{\mathbb{R}^{d}}x_{i}\varphi_{\varepsilon}(y)dy=s_{i}$ for some $i\in\{1,\dots,d\}$ and all $x\in\mathbb{R}^{d}$. Thus if $p\in\mathscr{P}_{1}(\mathbb{R}^{d})$ we get
@@
that is, polynomials in $\mathscr{P}_{1}(\mathbb{R}^{d})$ and $\mathscr{P}_{0}(\mathbb{R}^{d})$ are invariant to the convolution with even Schwartz convolution kernel. Thus, we can reach the same conclusions as in Theorem 2.12 with an identical proof. That is, convolution with an even convolution kernel is an isomorphism of $\mathscr{P}_{n}(\mathbb{R}^{d})$ onto itself for all $n\in\mathbb{N}_{0}$, and we can compute the inverse of the convolution similar to equations (7) and (8). Indeed, if $n$ is even the inverse of $p\in\mathscr{P}_{n}(\mathbb{R}^{d})\mapsto T_{\varepsilon}(p)=p*\varphi_{\varepsilon}\in C^{\infty}(\mathbb{R}^{d})$ with $\varphi\in\mathscr{S}_{M}(\mathbb{R}^{d})$ even is
@@ 3 Applications
In this section we present two particular applications of the results in Section 2. First we validate numerically Theorem 2.12, and then we apply the results of Theorem 2.18 to an application in the reconstruction of a sinusoidal function.
Figure 1: Representation of a numerical validation of Theorem 2.12. The left picture represents as a solid black line the original polynomial $p$ which is the $n=50$ order Taylor polynomial of the function $x\in\mathbb{R}\mapsto\sin(5x)+\sin(3x)$. It also represents as a blue solid line its convolution $p*\varphi_{\varepsilon}$ with convolution kernel $s\mapsto\varphi(x)=\operatorname{ind}_{(-1,1)}\exp(-1/(1-x^{2}))$ and parameter $\varepsilon=0.9$. The dashed red line represents $T_{\varepsilon}^{-1}(p*\varphi_{\varepsilon})$ computed using equation (7). The right picture represents the error between the original polynomial and the inverse of the convolution.
3.1 Validation of Theorem 2.12
In order to numerically validate Theorem 2.12, we present Figure 1. The polynomial is an $n=50$ Taylor polynomial of the function $x\in\mathbb{R}\mapsto\sin(3x)+\sin(5x)$. Due to the fact that, numerically speaking, the domain of the original polynomial is always bounded, it can be seen that at the ends of the domain the inverse of the convolution and the original polynomial differs the most. This is a fact on how convolution is computed numerically. Due to being the functions actually finite sequences, a zero padding is needed in order to be able to compute the convolution numerically. Then, once computed, we take back the same number of samples as the original sequence in order to compare it with the original function in its domain of definition. Nevertheless, it can be seen that, once the effect of the edges is negligible, that Figure 1 validates numerically Theorem 2.12, recovering $p$ from $p*\varphi_{\varepsilon}$ by applying Equation (7). Note that the used convolution kernel has compact support and zeros in its Fourier transform (cf. Theorem 2.18).
@@ 3.2 Applications to Signal Processing
We are going to consider the problem of deconvolution of an arbitrary function. Suppose that the convolution kernel is the Gaussian Kernel, $\varphi\in\mathscr{S}(\mathbb{R})$. Suppose then, that some function $f:\mathbb{R}\to\mathbb{R}$ has been convolved with $\varphi_{\varepsilon}$ for some $\varepsilon>0$. Suppose further that $f$ satisfies any condition of Theorem 2.18, Then, by letting $n\to\infty$ in Equation (7) we could recover $f$ (in the appropriate convergence).
The traditional approach to deconvolution consists on working on the Fourier transformed space. For example, under the appropriate conditions, for $\varphi$ the Gaussian Kernel and $f\in\mathscr{S}(\mathbb{R})$ it holds that $\widehat{f*\varphi}(\xi)=\widehat{f}\widehat{\varphi}(\xi)\Longrightarrow f=\mathcal{F}^{-1}\left(\widehat{f*\varphi}e^{\operatorname{id}^{2}}\right)$. Note this is ill-posed numerically speaking, since $x\mapsto e^{x^{2}}$ growths exponentially fast, and in particular, this function does not posses inverse Fourier transform. However, by using the results of Theorem 2.18, we can reconstruct $f$ directly from the samples of $f*\varphi$, without the need of passing to the frequency domain.
Consider then, the function $t\in\mathbb{R}\mapsto f(t)=\sin(5t)+\sin(3t)$ (we denote $t$ for time) as in Figure 1 and let $\varphi$ be the Gaussian convolution kernel. Since $f$ defines a regular tempered distribution, the assumptions of Theorem 2.18 hold.
Figure 2: Representation of the numerical validation of Theorem 2.18 under ideal conditions. Left picture represents in black the original function $t\mapsto f(t)=\sin(5t)+\sin(3t)$, in blue its convolution $T_{\varepsilon}(f)$ with the Gaussian Kernel $\varphi$ with $\varepsilon=0.55$, and as dashed red line its reconstruction $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$ using Equation (9) for $n=90$. Right picture represents in black the Fast Fourier Transform (FFT) of $f$, in green the FFT of $\varphi$, in blue the FFT of $T_{\varepsilon}(f)$ and in red the FFT of the reconstructed function $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$.
Note $T_{\varepsilon}(f)=f*\varphi_{\varepsilon}$ for $\varepsilon>0$ is the available data. In Figure 2 we can see the results of applying Theorem 2.18 to $T_{\varepsilon}(f)$. As it can be seen, the signal reconstruction is almost perfect. Theoretically speaking, we would need to let $n\to\infty$ for (distributional) convergence, but at $n\geq 100$ the numerical errors start to affect the method, and the recovery of the signal for $n=90$ is sufficiently good, numerically speaking, as can be seen from the time and frequency domains. Indeed, by looking at the spectra of $f$, $T_{\varepsilon}(f)$ and $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f))$, it is well validated that the method works, since we were capable of recovering (numerically speaking) the original function.
@@
Figure 3: Representation of the numerical validation of Theorem 2.18 under non ideal conditions. Top and bottom left pictures represent in black the original function $t\mapsto f(t)=\sin(5t)+\sin(3t)$, with a blue line $T_{\varepsilon}(f)$ with $\varepsilon=0.55$ and $\varphi$ the Gaussian Kernel, as a yellow line $T_{\varepsilon}(f)+\eta$, where $\eta$ is noise generated using a normal distribution with zero mean and variance 0, and with a red line $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)$ where $n=90$. Top right picture depicts as a black line the FFT of $f$, as a yellow line the FFT of $T_{\varepsilon}(f)+\eta$ and as a red line $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)$, i.e., the application of Equation (9) to $T_{\varepsilon}(f)+\eta$. The bottom right picture shows with a black line the FFT of $f$, with a green line the FFT of the used filter $t\mapsto h(t)=2\operatorname{sinc}(2t)$, with a yellow line the FFT of $T_{\varepsilon}(f)+\eta$ and with a red line the FFT of $T_{\varepsilon,n}^{-1}(T_{\varepsilon}(f)+\eta)*h$, that is, $g:=T_{\varepsilon}(f)+\eta$.
@@ 4 Conclusions
This work established a rigorous and computationally robust framework for the exact inversion of convolution operators with even Schwartz kernels. We first proved that these operators act as automorphisms on the vector space of finite-order polynomials, yielding an explicit, finite-sum inversion formula based solely on recursive forward convolutions. By extending this algebraic foundation to infinite-dimensional spaces, we derived a novel iterative formula for recovering functions and tempered distributions in their respective topologies. This approach provides a new alternative to traditional differential inversion of the Weierstrass transform, bypassing the numerical ill-posedness inherent in such operators. Numerical validation confirmed the feasibility of the method, demonstrating successful signal recovery for sinusoidal functions in both noiseless and low signal-to-noise ratio — less than unity — environments. These results confirm that the recursive convolution framework is not only theoretically exact but also highly resilient for practical signal processing and computational applications